What is the water solubility of Baco; (Ksp - 2.6 x 10-9 in a solution containing 040 M CO, (aq)? Answer: Check
Answer #1

[CO32-] = 0.4 M

At equilibrium:

BaCO3 <----> Ba2+ + CO32-

   s 0.4 + s

Ksp = [Ba2+][CO32-]

2.6*10^-9=(s)*(0.4+ s)

Since Ksp is small, s can be ignored as compared to 0.4

Above expression thus becomes:

2.6*10^-9=(s)*(0.4)

2.6*10^-9= (s) * 0.4

s = 6.5*10^-9 M

Answer: 6.5*10^-9 M